# Generalised rank-constrained approximations of Hilbert–Schmidt operators on separable Hilbert spaces and applications

### v1

@@ 3 Optimal approximations for Hilbert–Schmidt operators on separable Hilbert spaces
We are now ready to formulate (3) and (4) with the minimality property (8) in infinite dimensions. In Section 3.1, we give necessary and sufficient conditions for a solution to exist and describe the solution. An interpretation of the minimality property is also given, as well as the finite-dimensional modification of [20, Theorem 2.1]. When no solution exists, i.e. when the minimum Hilbert–Schmidt approximation error is not attained, one can still explicitly construct arbitrarily good approximations, as is shown in Section 3.2. When a solution does exist, the corresponding optimal Hilbert–Schmidt approximation error is computed in Section 3.3.
@@ 3.1 Problem formulation and solution
$\mathcal{X}_{r}$$\coloneqq\left\{X:\mathcal{H}_{2}\rightarrow\mathcal{H}_{3},\ \operatorname{dom}{X}\supset\operatorname{dom}{C^{\dagger}},\right.$
$\left.\qquad\dim\operatorname{ran}{X}\leq r,BXC\in L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})\right\},$
$\mathcal{Y}_{r}$$\coloneqq\left\{Y\in\mathcal{B}_{00,r}(\mathcal{H}_{1},\mathcal{H}_{4}):\ \forall\tilde{Y}\in\mathcal{B}_{00,r}(\mathcal{H}_{1},\mathcal{H}_{4})\right.$
$\left.\qquad\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-Y\rVert_{HS}\leq\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-\tilde{Y}\rVert_{HS},\right.$
$\left.\qquad\operatorname{ran}{Y}\subset\operatorname{dom}{B^{\dagger}}\right\}.$
@@
$\min\{\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-Y\rVert_{HS}:\ Y\in\mathcal{B}_{00,r}(\mathcal{H}_{1},\mathcal{H}_{4})\},$ (9)
that also map into $\operatorname{dom}{B^{\dagger}}$. Hence Proposition A.7, $\mathcal{Y}_{r}$ consists of all rank-$r$ truncated SVDs of $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$, as defined in (38), that map into $\operatorname{dom}{B^{\dagger}}$.
@@
$\min\{\lVert M-BXC\rVert_{HS},\ X\in\mathcal{X}_{r}\},$ (10)
@@
$\hat{X}:\operatorname{dom}{C^{\dagger}}\subset\mathcal{H}_{2}\rightarrow\mathcal{H}_{3},\quad\hat{X}=B^{\dagger}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}C^{\dagger},$ (11)
@@
This solution is unique, in the sense that $X=\tilde{X}$ on $\operatorname{dom}{C^{\dagger}}$ for solutions $X,\tilde{X}$ of the form (11) satisfying (12), if and only if one of the following conditions holds:
@@
By Proposition A.5 item (iv),
@@
$\mathcal{S}_{r}$$\coloneqq\left\{X\in\mathcal{X}_{r}:\ \lVert M-BXC\rVert_{HS}\leq\lVert M-B\tilde{X}C\rVert_{HS}\ \forall\tilde{X}\in\mathcal{X}_{r}\right\},$
$\mathcal{S}_{r,*}$$\coloneqq\left\{X\in\mathcal{S}_{r}:\ X=P_{\ker{B}^{\perp}}XP_{\overline{\operatorname{ran}{C}}}\text{ on }\operatorname{dom}{X}\right\},$
Thus, $\mathcal{S}_{r}$ is the set of solutions to the reduced rank problem (10). Further, $\mathcal{S}_{r,*}$ is the set of solutions to (10) that also satisfy (12). We must show, under the existence condition $\mathcal{Y}_{r}\not=\emptyset$, that $\hat{X}$ defined in (11) lies in $\mathcal{S}_{r,*}$ and, under the uniqueness condition (a) or (b), that $X=\tilde{X}$ on $\operatorname{dom}{C^{\dagger}}$ for any $X,\tilde{X}\in\mathcal{S}_{r,*}$. To this end, we show that solutions to (10) and (12) can be constructed by a solution to (9) which maps into $\operatorname{dom}{B^{\dagger}}$ and vice versa, thereby drawing a connection between the sets $\mathcal{S}_{r,*}$ and $\mathcal{Y}_{r}$. The following lemma is crucial for this purpose.
@@
Let $M$, $B$, $C$ and $\mathcal{X}_{r}$ be as in Theorem 3.2 and $c\coloneqq\lVert M\rVert_{HS}^{2}-\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}\rVert_{HS}^{2}$. For any $X\in\mathcal{X}_{r}$,
$\lVert M-BXC\rVert_{HS}^{2}$ $=\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-BXC\rVert_{HS}^{2}+c.$ (13)
@@
$\langle P_{\overline{\operatorname{ran}{B}}}S,(I-P_{\overline{\operatorname{ran}{B}}})T\rangle_{HS}=\sum_{i}\langle P_{\overline{\operatorname{ran}{B}}}Se_{i},(I-P_{\overline{\operatorname{ran}{B}}})Te_{i}\rangle=0,\quad T,S\in\mathcal{B}(\mathcal{K},\mathcal{H}_{4}),$
$\langle SP_{\ker{C}^{\perp}},T(I-P_{\ker{C}^{\perp}})\rangle_{HS}=\langle(I-P_{\ker{C}^{\perp}})T^{*},P_{\ker{C}^{\perp}}S^{*}\rangle_{HS}=0,\quad T,S\in\mathcal{B}(\mathcal{H}_{1},\mathcal{K}).$
Furthermore, the Moore–Penrose equations (34), (36) and (37) in Proposition A.5 state that $P_{\overline{\operatorname{ran}{B}}}B=BB^{\dagger}B=B$ and $CP_{\ker{C}^{\perp}}=CC^{\dagger}C=C$. Therefore, for any $X\in\mathcal{X}_{r}$, we have $P_{\overline{\operatorname{ran}{B}}}BXCP_{\ker{C}^{\perp}}=BXC\in L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$ and,
$\lVert M-BXC\rVert_{HS}^{2}$ $=$ $\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-BXC\rVert_{HS}^{2}+\lVert(I-P_{\overline{\operatorname{ran}{B}}})MP_{\ker{C}^{\perp}}\rVert_{HS}^{2}$
$+\lVert P_{\overline{\operatorname{ran}{B}}}M(I-P_{\ker{C}^{\perp}})\rVert_{HS}^{2}+\lVert(I-P_{\overline{\operatorname{ran}{B}}})M(I-P_{\ker{C}^{\perp}})\rVert_{HS}^{2}$
$=$ $\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-BXC\rVert_{HS}^{2}+c.$
@@
Let $M,B,C$ and $\mathcal{X}_{r}$ be as in Theorem 3.2. Then
@@
    For $Y\in\mathcal{Y}_{r}$ it holds that ${B^{\dagger}YC^{\dagger}}$ is a well-defined linear operator on $\operatorname{dom}{C^{\dagger}}$, which satisfies $B^{\dagger}YC^{\dagger}\in\mathcal{S}_{r}$ and $BB^{\dagger}YC^{\dagger}C=Y$,
@@
    For $X\in\mathcal{X}_{r}$, $BXC\in\mathcal{Y}_{r}$ if and only if $X\in\mathcal{S}_{r}$,
@@
    For $Y\in\mathcal{Y}_{r}$ it holds that $B^{\dagger}YC^{\dagger}\in\mathcal{S}_{r,*}$,
@@
If we define the map $j:\mathcal{X}_{r}\rightarrow\mathcal{B}_{00,r}(\mathcal{H}_{1},\mathcal{H}_{4})$, $X\mapsto BXC$, then the “if” part of (ii) in Lemma 3.7 implies that $j(\mathcal{S}_{r})\subset\mathcal{Y}_{r}$ and the “only if” part that $j(\mathcal{X}_{r}\backslash\mathcal{S}_{r})\subset\mathcal{B}_{00,r}(\mathcal{H}_{1},\mathcal{H}_{4})\backslash\mathcal{Y}_{r}$. Item (i) states that ${\left.\kern-1.2ptj\vphantom{\big{|}}\right|_{\mathcal{S}_{r}}}$ maps onto $\mathcal{Y}_{r}$, hence has a right inverse, and that a right inverse is given by the map $Y\mapsto B^{\dagger}YC^{\dagger}$ on $\operatorname{dom}{C^{\dagger}}$. By item (iii), the image of this right inverse maps into $\mathcal{S}_{r,*}$. By item (iv), every $X\in\mathcal{S}_{r,*}$ agrees with $B^{\dagger}YC^{\dagger}$ on $\operatorname{dom}{C^{\dagger}}$, for some $Y\in\mathcal{Y}_{r}$.
@@
In particular, $B(B^{\dagger}YC^{\dagger})C$ has rank at most $r$ and is hence Hilbert–Schmidt. Thus, $B^{\dagger}YC^{\dagger}\in\mathcal{X}_{r}$. Finally, $B^{\dagger}YC^{\dagger}\in\mathcal{S}_{r}$, since for any $\tilde{X}\in\mathcal{X}_{r}$ it holds by applying (13) twice, (14) and the definition of $Y$,
$\lVert M-B\tilde{X}C\rVert_{HS}^{2}$ $=\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-B\tilde{X}C\rVert_{HS}^{2}+c\geq\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-Y\rVert_{HS}^{2}+c$
$=\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-BB^{\dagger}YC^{\dagger}C\rVert_{HS}^{2}+c=\lVert M-B(B^{\dagger}YC^{\dagger})C\rVert_{HS}^{2}.$
@@
$\lVert M-B\tilde{X}C\rVert_{HS}^{2}$ $=\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-B\tilde{X}C\rVert_{HS}^{2}+c$
$\geq\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-B{X}C\rVert_{HS}^{2}+c=\lVert M-B{X}C\rVert_{HS}^{2}.$
@@
$\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-\tilde{Y}\rVert_{HS}^{2}$ $\geq\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-Y\rVert_{HS}^{2}=\lVert M-B(B^{\dagger}YC^{\dagger})C\rVert_{HS}^{2}-c$
$\geq\lVert M-BXC\rVert_{HS}^{2}-c=\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-BXC\rVert_{HS},$
@@
(iii) Let $Y\in\mathcal{Y}_{r}$. By item (i), $B^{\dagger}YC^{\dagger}\in\mathcal{S}_{r}$. By (35),
@@
(iv) Let $X\in\mathcal{S}_{r,*}$, so $Y\coloneqq BXC\in\mathcal{Y}_{r}$ by item (ii). By definition of $\mathcal{S}_{r,*}$ it holds that $X=P_{\ker{B}^{\perp}}XP_{\overline{\operatorname{ran}{C}}}$ on $\operatorname{dom}{C^{\dagger}}\subset\operatorname{dom}{X}$. Hence, using (36) and (37), $X=B^{\dagger}(BXC)C^{\dagger}=B^{\dagger}YC^{\dagger}$ on $\operatorname{dom}{C^{\dagger}}$.
@@
It remains to show that all elements of $S_{r,*}$ agree on $\operatorname{dom}{C^{\dagger}}$ under conditions (a) or (b). By the construction of the SVD in Section A.3, $\mathcal{Y}_{r}$ contains at most one element if and only if (a) or (b) holds. In this case, by item (iv) of Lemma 3.7, all elements of $\mathcal{S}_{r,*}$ agree on $\operatorname{dom}{C^{\dagger}}$. This concludes the proof of uniqueness. ∎
@@
Let $B\in\mathbb{C}^{m\times p},C\in\mathbb{C}^{q\times n},M\in\mathbb{C}^{m\times n}.$ Then
@@
$\min\{\lVert M-BXC\rVert_{F},\ X\in\mathbb{C}^{p\times q},\operatorname{rank}\left({X}\right)\leq r\}.$
Furthermore, it satisfies
@@
This solution is unique if and only if one of the following conditions holds,
@@
We associate a matrix $A\in\mathbb{C}^{t\times s}$ with the linear operator $L_{A}\in\mathcal{B}(\mathbb{C}^{s},\mathbb{C}^{t})$ that left multiplies $x\in\mathbb{C}^{s}$ with $A$, i.e $L_{A}x=Ax$. Hence $A$ is the expression of $L_{A}$ in the orthogonal standard bases of $\mathbb{C}^{s}$ and $\mathbb{C}^{t}$, so we have the equality of norms $\lVert L_{A}\rVert_{HS}=\lVert A\rVert_{F}$. Furthermore, one may check that $L_{A}^{\dagger}=L_{A^{\dagger}}$, $(L_{A})_{r}=L_{(A)_{r}}$, $P_{\overline{\operatorname{ran}{L_{A}}}}=P_{\operatorname{ran}{L_{A}}}=P_{\operatorname{ran}{A}}=L_{P_{A,L}}$ and $P_{\ker{L_{A}}^{\perp}}=P_{\ker{A}^{\perp}}=L_{P_{A,R}}$. As all finite-dimensional linear maps are Hilbert–Schmidt operators and have closed range, Remark 3.5 implies that we can apply Theorem 3.2 to $L_{M},L_{B},L_{C}$ with $\mathcal{X}_{r}=\mathcal{B}_{00,r}(\mathbb{C}^{q},\mathbb{C}^{p})\simeq\{X\in\mathbb{C}^{p\times q},\operatorname{rank}\left({X}\right)\leq r\}$ and where $\mathcal{Y}_{r}$ consists of the truncated rank-$r$ SVDs of $P_{{\operatorname{ran}{B}}}L_{M}P_{\ker{C}^{\perp}}$. The result now follows from the conclusions of Theorem 3.2 by expressing the linear operators again as matrices. ∎
@@
Let us decompose $\mathcal{H}_{3}=\ker{B}^{\perp}\oplus\ker{B}$ and $\operatorname{dom}{X}=\mathcal{D}\oplus\operatorname{ran}{C}^{\perp}$ We define,
$X_{11}$ $\coloneqq{\left.\kern-1.2pt\left(P_{\ker{B}^{\perp}}XP_{\overline{\operatorname{ran}{C}}}\right)\vphantom{\big{|}}\right|_{\mathcal{D}}}:\mathcal{D}\subset\overline{\operatorname{ran}{C}}\rightarrow\ker{B}^{\perp}$
$X_{12}$ $\coloneqq{\left.\kern-1.2pt\left(P_{\ker{B}^{\perp}}XP_{\operatorname{ran}{C}^{\perp}}\right)\vphantom{\big{|}}\right|_{\operatorname{ran}{C}^{\perp}}}:\operatorname{ran}{C}^{\perp}\rightarrow\ker{B}^{\perp}$
$X_{21}$ $\coloneqq{\left.\kern-1.2pt\left(P_{\ker{B}}XP_{\overline{\operatorname{ran}{C}}}\right)\vphantom{\big{|}}\right|_{\mathcal{D}}}:\mathcal{D}\subset\overline{\operatorname{ran}{C}}\rightarrow\ker{B}$

### v3

@@ 3 Optimal approximations for Hilbert–Schmidt operators on separable Hilbert spaces
We are now ready to formulate (3) and (4) with the minimality property (8) in infinite dimensions. In Section 3.1, we give necessary and sufficient conditions for a solution to exist and describe the solution. An interpretation of the minimality property is also given, as well as the finite-dimensional modification of [20, Theorem 2.1]. When no solution exists, i.e. when the minimum Hilbert–Schmidt approximation error is not attained, one can still explicitly construct arbitrarily good approximations, as is shown in Section 3.2. When a solution does exist, the corresponding optimal Hilbert–Schmidt approximation error is computed in Section 3.3. The theory below involves the use of unbounded linear operators, SVDs of compact operators and Moore–Penrose inverses, which are recalled in Appendix A.
@@ 3.1 Problem formulation and solution
$\mathcal{X}_{r}$$\coloneqq\left\{X:\mathcal{H}_{2}\rightarrow\mathcal{H}_{3}:\ \operatorname{dom}{X}\supset\operatorname{dom}{C^{\dagger}},\dim\operatorname{ran}{X}\leq r,BXC\in L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})\right\},$
$\mathcal{Y}_{r}$$\coloneqq\left\{Y\in\mathcal{B}_{00,r}(\mathcal{H}_{1},\mathcal{H}_{4}):\ \operatorname{ran}{Y}\subset\operatorname{dom}{B^{\dagger}},\right.$
$\left.\qquad\forall\tilde{Y}\in\mathcal{B}_{00,r}(\mathcal{H}_{1},\mathcal{H}_{4}),\ \lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-Y\rVert_{L_{2}}\leq\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-\tilde{Y}\rVert_{L_{2}}\right\}.$
@@
$\min\{\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-Y\rVert_{L_{2}}:\ Y\in\mathcal{B}_{00,r}(\mathcal{H}_{1},\mathcal{H}_{4})\},$ (9)
that also map into $\operatorname{dom}{B^{\dagger}}$. Hence by Proposition A.7, $\mathcal{Y}_{r}$ consists of all rank-$r$ truncated SVDs of $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$, as defined in (40), that map into $\operatorname{dom}{B^{\dagger}}$.
@@
$\min\{\lVert M-BXC\rVert_{L_{2}}:\ X\in\mathcal{X}_{r}\}$ (10)
@@
$\hat{X}:\operatorname{dom}{C^{\dagger}}\subset\mathcal{H}_{2}\rightarrow\mathcal{H}_{3},\quad\hat{X}=B^{\dagger}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}C^{\dagger}.$ (11)
@@
This solution is unique, in the sense that $X=\tilde{X}$ on $\operatorname{dom}{C^{\dagger}}$ for solutions $X,\tilde{X}$ of (10) satisfying (12), if and only if one of the following conditions holds:
@@
By Proposition A.5 item (iv), the following hold:
@@
$\mathcal{S}_{r}$$\coloneqq\left\{X\in\mathcal{X}_{r}:\ \forall\tilde{X}\in\mathcal{X}_{r},\ \lVert M-BXC\rVert_{L_{2}}\leq\lVert M-B\tilde{X}C\rVert_{L_{2}}\right\},$
$\mathcal{S}_{r,*}$$\coloneqq\left\{X\in\mathcal{S}_{r}:\ X=P_{\ker{B}^{\perp}}XP_{\overline{\operatorname{ran}{C}}}\text{ on }\operatorname{dom}{X}\right\}.$
Thus, $\mathcal{S}_{r}$ is the set of solutions to the reduced rank problem (10). Further, $\mathcal{S}_{r,*}$ is the set of solutions to (10) that also satisfy (12). We must show, under the existence condition $\mathcal{Y}_{r}\not=\emptyset$, that $\hat{X}$ defined in (11) lies in $\mathcal{S}_{r,*}$, and that $X=\tilde{X}$ on $\operatorname{dom}{C^{\dagger}}$ for any $X,\tilde{X}\in\mathcal{S}_{r,*}$ if and only if the uniqueness condition (a) or (b) holds. To this end, we show that solutions to (10) and (12) can be constructed by a solution to (9) which maps into $\operatorname{dom}{B^{\dagger}}$ and vice versa, thereby drawing a connection between the sets $\mathcal{S}_{r,*}$ and $\mathcal{Y}_{r}$. The following lemma is crucial for this purpose.
@@
Let $M$, $B$, $C$ and $\mathcal{X}_{r}$ be as in Theorem 3.2 and $c\coloneqq\lVert M\rVert_{L_{2}}^{2}-\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}\rVert_{L_{2}}^{2}$. For any $X\in\mathcal{X}_{r}$,
$\lVert M-BXC\rVert_{L_{2}}^{2}$ $=\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-BXC\rVert_{L_{2}}^{2}+c.$ (13)
@@
$\langle P_{\overline{\operatorname{ran}{B}}}S,(I-P_{\overline{\operatorname{ran}{B}}})T\rangle_{L_{2}}=\sum_{i}\langle P_{\overline{\operatorname{ran}{B}}}Se_{i},(I-P_{\overline{\operatorname{ran}{B}}})Te_{i}\rangle=0,\quad T,S\in L_{2}(\mathcal{K},\mathcal{H}_{4}),$
$\langle SP_{\ker{C}^{\perp}},T(I-P_{\ker{C}^{\perp}})\rangle_{L_{2}}=\langle(I-P_{\ker{C}^{\perp}})T^{*},P_{\ker{C}^{\perp}}S^{*}\rangle_{L_{2}}=0,\quad T,S\in L_{2}(\mathcal{H}_{1},\mathcal{K}).$
Furthermore, the Moore–Penrose equations (36), (38) and (39) in Proposition A.5 state that $P_{\overline{\operatorname{ran}{B}}}B=BB^{\dagger}B=B$ and $CP_{\ker{C}^{\perp}}=CC^{\dagger}C=C$. Therefore, for any $X\in\mathcal{X}_{r}$, we have $P_{\overline{\operatorname{ran}{B}}}BXCP_{\ker{C}^{\perp}}=BXC\in L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$, and
$\lVert M-BXC\rVert_{L_{2}}^{2}$ $=$ $\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-BXC\rVert_{L_{2}}^{2}+\lVert(I-P_{\overline{\operatorname{ran}{B}}})MP_{\ker{C}^{\perp}}\rVert_{L_{2}}^{2}$
$+\lVert P_{\overline{\operatorname{ran}{B}}}M(I-P_{\ker{C}^{\perp}})\rVert_{L_{2}}^{2}+\lVert(I-P_{\overline{\operatorname{ran}{B}}})M(I-P_{\ker{C}^{\perp}})\rVert_{L_{2}}^{2}$
$=$ $\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-BXC\rVert_{L_{2}}^{2}+c.$
@@
Let $M$, $B$, $C$ and $\mathcal{X}_{r}$ be as in Theorem 3.2. Then the following statements hold.
@@
    For $Y\in\mathcal{Y}_{r}$ it holds that ${B^{\dagger}YC^{\dagger}}$ is a well-defined linear operator on $\operatorname{dom}{C^{\dagger}}$, which satisfies $B^{\dagger}YC^{\dagger}\in\mathcal{S}_{r}$ and $BB^{\dagger}YC^{\dagger}C=Y$.
@@
    For $X\in\mathcal{X}_{r}$, $BXC\in\mathcal{Y}_{r}$ if and only if $X\in\mathcal{S}_{r}$.
@@
    For $Y\in\mathcal{Y}_{r}$ it holds that $B^{\dagger}YC^{\dagger}\in\mathcal{S}_{r,*}$.
@@
If we define the map $j:\mathcal{X}_{r}\rightarrow\mathcal{B}_{00,r}(\mathcal{H}_{1},\mathcal{H}_{4})$, $X\mapsto BXC$, then the “if” part of (ii) in Lemma 3.7 implies that $j(\mathcal{S}_{r})\subset\mathcal{Y}_{r}$ and the “only if” part implies that $j(\mathcal{X}_{r}\backslash\mathcal{S}_{r})\subset\mathcal{B}_{00,r}(\mathcal{H}_{1},\mathcal{H}_{4})\backslash\mathcal{Y}_{r}$. Item (i) states that ${\left.\kern-1.2ptj\vphantom{\big{|}}\right|_{\mathcal{S}_{r}}}$ maps onto $\mathcal{Y}_{r}$, hence has a right inverse, and that a right inverse is given by the map $Y\mapsto B^{\dagger}YC^{\dagger}$ on $\operatorname{dom}{C^{\dagger}}$. By item (iii), the image of this right inverse maps into $\mathcal{S}_{r,*}$. By item (iv), every $X\in\mathcal{S}_{r,*}$ agrees with $B^{\dagger}YC^{\dagger}$ on $\operatorname{dom}{C^{\dagger}}$, for some $Y\in\mathcal{Y}_{r}$.
@@
In particular, $B(B^{\dagger}YC^{\dagger})C$ has rank at most $r$ and is hence Hilbert–Schmidt. Thus, $B^{\dagger}YC^{\dagger}\in\mathcal{X}_{r}$. Finally, $B^{\dagger}YC^{\dagger}\in\mathcal{S}_{r}$, since for any $\tilde{X}\in\mathcal{X}_{r}$, applying (13) twice, (14) and the definition of $Y$ yields
$\lVert M-B\tilde{X}C\rVert_{L_{2}}^{2}$ $=\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-B\tilde{X}C\rVert_{L_{2}}^{2}+c\geq\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-Y\rVert_{L_{2}}^{2}+c$
$=\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-BB^{\dagger}YC^{\dagger}C\rVert_{L_{2}}^{2}+c=\lVert M-B(B^{\dagger}YC^{\dagger})C\rVert_{L_{2}}^{2}.$
@@
$\lVert M-B\tilde{X}C\rVert_{L_{2}}^{2}$ $=\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-B\tilde{X}C\rVert_{L_{2}}^{2}+c$
$\geq\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-B{X}C\rVert_{L_{2}}^{2}+c=\lVert M-B{X}C\rVert_{L_{2}}^{2}.$
@@
$\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-\tilde{Y}\rVert_{L_{2}}^{2}$ $\geq\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-Y\rVert_{L_{2}}^{2}=\lVert M-B(B^{\dagger}YC^{\dagger})C\rVert_{L_{2}}^{2}-c$
$\geq\lVert M-BXC\rVert_{L_{2}}^{2}-c=\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-BXC\rVert_{L_{2}},$
@@
(iii) Let $Y\in\mathcal{Y}_{r}$. By item (i), $B^{\dagger}YC^{\dagger}\in\mathcal{S}_{r}$. By (37), (38) and (39),
@@
(iv) Let $X\in\mathcal{S}_{r,*}\subset\mathcal{S}_{r}$, so $Y\coloneqq BXC\in\mathcal{Y}_{r}$ by item (ii). By definition of $\mathcal{S}_{r,*}$ it holds that $X=P_{\ker{B}^{\perp}}XP_{\overline{\operatorname{ran}{C}}}$ on $\operatorname{dom}{C^{\dagger}}\subset\operatorname{dom}{X}$. Hence, using (38) and (39), $X=B^{\dagger}(BXC)C^{\dagger}=B^{\dagger}YC^{\dagger}$ on $\operatorname{dom}{C^{\dagger}}$.
@@
It remains to show that all elements of $S_{r,*}$ agree on $\operatorname{dom}{C^{\dagger}}$ if and only if condition (a) or (b) holds. By the construction of the SVD in Section A.3, $\mathcal{Y}_{r}$ contains at most one element if and only if (a) or (b) holds. If $\mathcal{Y}_{r}$ contains at most one element, then by item (iv) of Lemma 3.7, all elements of $\mathcal{S}_{r,*}$ agree on $\operatorname{dom}{C^{\dagger}}$. Conversely, if all elements of $S_{r,*}$ agree on $\operatorname{dom}{C^{\dagger}}$, and if $Y_{1},Y_{2}\in\mathcal{Y}_{r}$, then $B^{\dagger}Y_{1}C^{\dagger}$ and $B^{\dagger}Y_{2}C^{\dagger}$ agree on $\operatorname{dom}{C}^{\dagger}$ by Lemma 3.7(iii). Hence by Lemma 3.7(i), $Y_{1}=B(B^{\dagger}Y_{1}C^{\dagger})C=B(B^{\dagger}Y_{2}C^{\dagger})C=Y_{2}$, so that $\mathcal{Y}_{r}$ has at most one element. This concludes the proof of uniqueness. ∎
@@
Let $B\in\mathbb{C}^{m\times p}$, $C\in\mathbb{C}^{q\times n}$ and $M\in\mathbb{C}^{m\times n}.$ Then
@@
$\min\{\lVert M-BXC\rVert_{F}:\ X\in\mathbb{C}^{p\times q},\operatorname{rank}\left({X}\right)\leq r\}.$
Furthermore, $\hat{X}$ satisfies
@@
This solution is the unique solution satisfying (15) if and only if one of the following conditions holds:
@@
We associate a matrix $A\in\mathbb{C}^{t\times s}$ with the linear operator $L_{A}\in\mathcal{B}(\mathbb{C}^{s},\mathbb{C}^{t})$ that left multiplies $x\in\mathbb{C}^{s}$ with $A$, i.e $L_{A}x=Ax$. Hence $A$ is the expression of $L_{A}$ in the orthogonal standard bases of $\mathbb{C}^{s}$ and $\mathbb{C}^{t}$, so we have the equality of norms $\lVert L_{A}\rVert_{L_{2}}=\lVert A\rVert_{F}$. Furthermore, one may check that $L_{A}^{\dagger}=L_{A^{\dagger}}$, $(L_{A})_{r}=L_{(A)_{r}}$, $P_{\overline{\operatorname{ran}{L_{A}}}}=P_{\operatorname{ran}{L_{A}}}=P_{\operatorname{ran}{A}}=L_{P_{A,L}}$ and $P_{\ker{L_{A}}^{\perp}}=P_{\ker{A}^{\perp}}=L_{P_{A,R}}$. As all finite-dimensional linear maps are Hilbert–Schmidt operators and have closed range, Remark 3.5 implies that we can apply Theorem 3.2 to $L_{M},L_{B},L_{C}$ with $\mathcal{X}_{r}=\mathcal{B}_{00,r}(\mathbb{C}^{q},\mathbb{C}^{p})\simeq\{X\in\mathbb{C}^{p\times q},\operatorname{rank}\left({X}\right)\leq r\}$ and where $\mathcal{Y}_{r}$ consists of the truncated rank-$r$ SVDs of $P_{{\operatorname{ran}{B}}}L_{M}P_{\ker{C}^{\perp}}$. The result now follows from the conclusions of Theorem 3.2 by expressing the linear operators again as matrices. ∎
@@
Let us decompose $\mathcal{H}_{3}=\ker{B}^{\perp}\oplus\ker{B}$ and $\operatorname{dom}{X}=\mathcal{D}\oplus\operatorname{ran}{C}^{\perp}$ We define
$X_{11}$ $\coloneqq{\left.\kern-1.2pt\left(P_{\ker{B}^{\perp}}XP_{\overline{\operatorname{ran}{C}}}\right)\vphantom{\big{|}}\right|_{\mathcal{D}}}:\mathcal{D}\subset\overline{\operatorname{ran}{C}}\rightarrow\ker{B}^{\perp},$
$X_{12}$ $\coloneqq{\left.\kern-1.2pt\left(P_{\ker{B}^{\perp}}XP_{\operatorname{ran}{C}^{\perp}}\right)\vphantom{\big{|}}\right|_{\operatorname{ran}{C}^{\perp}}}:\operatorname{ran}{C}^{\perp}\rightarrow\ker{B}^{\perp},$
$X_{21}$ $\coloneqq{\left.\kern-1.2pt\left(P_{\ker{B}}XP_{\overline{\operatorname{ran}{C}}}\right)\vphantom{\big{|}}\right|_{\mathcal{D}}}:\mathcal{D}\subset\overline{\operatorname{ran}{C}}\rightarrow\ker{B},$